The function W(d)=0.112d approximatesthe amounts, in centimeters, of waterthat results from d centimeters of snow melting. Find the amount of water that results in snow melting from depths of 12 cm, 49cm, and 87cm
Answer by nyc_function(2741)

12 cm
W(12) = 0.112*(12) = 1.344 cm
49 cm
W(49) = 0.112*(49) = 5.488
87 cm
W(87) = 0.112*(87) = 9.744